OpenStudy (moongazer):

5^x = z , then whts 5^(x+3) is the answer 125z?

OpenStudy (anonymous):

yes it is 125z

OpenStudy (moongazer):

thanks!

OpenStudy (anonymous):

see 5^(x+3) = 5^x *5^3 because a^m * a^n = a^m+n = z*125 = 125z

OpenStudy (anonymous):

if u r satisfied, pls click GOOD ANSWER button :})

OpenStudy (moongazer):

I already know how to do it i just need a confirmation but i'll still give you a medal if your click the good answer button on beside my name :)

OpenStudy (anonymous):

did it ( :})

OpenStudy (moongazer):

^_^